# Pretrained Hybrids with MAD Skills

### v1

@@ Appendix A Ablations
In-context Recall 0.0068 0.0081
Fuzzy In-context Recall 4.1764 4.1729
Noisy In-context Recall 4.1628 4.1614
@@ Appendix C Additional Pathfinder Results
Appendix D Hyperparameters
D.1 Fine-Tuning Pretrained Hybrids
@@
Penn Treebank completions synthetic. For model weights, we use the AdamW [24] optimizer with a linear learning rate schedule with an initial learning rate of $55e-$ 5. For mixture weights, we use the AdamW [24] optimizer with a linear learning rate schedule with an initial learning rate of $0.005$ and use alternating updates.
@@
Fine-tuning on language tasks. For model weights, we use the AdamW [24] optimizer with a linear learning rate schedule with an initial learning rate of $55e-$ 5. For mixture weights, we use the AdamW [24] optimizer with a linear learning rate schedule with an initial learning rate of $0.005$ and use simultaneous updates.
D.2 Training Hybrids from Scratch
@@
Existing hybrids were trained with a hyperparameter grid search over the space $[1e-4,5e-4,1e-3]$ for learning rate and $[0.0,0.1]$ for weight decay, similar to the procedure in MAD [34].
Manticore is trained in two stages. In the first stage, we train the model and architecture weights in the alternating schedule utilized in DARTS [23]. In this stage, we perform a hyperparameter grid search of the space $[1e-4,5e-4,1e-3]$ for model weight learning rate, $[1e-4,1e-4]$ for architecture weight learning rate, and $[0.1]$ for weight decay. In the second stage, the architecture weights are frozen and we train only the model weights using the best learning rate found in the first stage.
@@
  • ListOps. We trained all models with 5000 steps. The hyperparameter for GPT-Neo is 8 heads, 6 layers, 512 as the embedding dimension, and 2048 as FFN dimension. The hyperparameter for Mamba is 12 layers, with 512 as the model dimension. The vocab size is 18.
  • IMDb. We trained all models with 25 epochs and batch size 32.The hyperparameter for GPT-Neo is 8 heads, 6 layers, 512 as the embedding dimension, and 2048 as FFN dimension. The hyperparameter for Mamba is 12 layers, with 512 as the model dimension. The vocab size is 129.
  • CIFAR10. We trained all models with 10 epochs. The hyperparameter for GPT-Neo is 4 heads, 3 layers, 64 as the embedding dimension, and 128 as FFN dimension. The hyperparameter for Mamba is 6 layers, with 64 as the model dimension. The vocab size is 256, which is the pixel value range of the grayscale image.
  • Pathfinder32. We trained all models with 10 epochs. The hyperparameter for GPT-Neo is 8 heads, 4 layers, 128 as the embedding dimension, and 128 as FFN dimension. The hyperparameter for Mamba is 8 layers, with 128 as the model dimension. The vocab size is 256, which is the pixel value range of the grayscale image.
@@
We use two blocks each from GPT-Neo and Mamba, each with a model dimension of 128. We train for 200 epochs and select the best performance during training, as all of the models overfit across the board. We use the AdamW [24] optimizer with a linear learning rate schedule with an initial learning rate of $55e-$ 5.
D.3 Programming Hybrids
@@
MAD tasks as a search proxy. For model weights, we use the AdamW [24] optimizer with a linear learning rate schedule with an initial learning rate of $55e-$ 5. For mixture weights, we use the AdamW [24] optimizer with a linear learning rate schedule with an initial learning rate of $0.01$ and use simultaneous updates. For search on the MAD tasks, we train scaled-down versions of GPT-Neo and Mamba each with four blocks, model dimensions of 128, and no projectors.
D.4 Pretraining Projectors
For all non-frozen weights (i.e., projectors, mixture weights, embeddings, and the LM head), we use the AdamW [24] optimizer with a linear learning rate schedule with an initial learning rate of $55e-$ 5.
Appendix E Data and MAD Task Parameters
@@
Natural language tasks. We evaluate the ability to fine-tune Manticore on natural language datasets. Specifically, we evaluate on Penn Treebank [26], the Alpaca instruction tuning dataset [37], and an i.i.d. split of the ELI5 training set [10]. Additionally, we use one billion tokens sampled from the FineWeb dataset [30] to pretrain our projector weights.
Appendix F A Call for Action & Community Recommendations
@@
We found that standard tools such as Huggingface and PyTorch were insufficient to cleanly access intermediate activations across several model implementations. This could be resolved by adopting standard implementations or structures for LMs that share the common block structure that we describe in Section 3.1.1. Instead, our solution was to fork implementations of several Huggingface models, which is time-consuming, error-prone, and non-scalable. A solution to this problem would enable and encourage further research on pretrained hybrid models, which in turn helps to democratize LM research.
@@ Appendix G Limitations
Appendix H Compute Resources
@@
Appendix I Broader Impacts and Safeguards
We acknowledge the possibility of misuse with respect to any form of LM research. In our work, among other things, we enable the creation of pretrained hybrid models from existing pretrained models. This has potentially positive and negative social impacts for the community. As a positive potential social impact, we enable the community to much more easily create their own hybrid models of various sizes without large scale pretraining—this has as much potential for positive impact in that these models can be used for good. On the other hand, the ability to create large pretrained hybrids, potentially with custom sets of skills, has the potential to open the door to misuse. To safeguard against such things, we will include appropriate licenses and rules for usage when we ultimately deploy a Python package for the community to more broadly use our framework.

### v2

@@ Appendix A Ablations
Context Recall 0.0068 0.0081
Fuzzy Recall 4.1764 4.1729
Noisy Recall 4.1628 4.1614
@@
Table 5: A comparison of non-discretized vs. discretized Manticore.
Amount of projector pretraining. Finally, we ablate over the amount of projector pretraining. We re-ran our $\alpha$ sweep on our PTB completions synthetic with different amounts of projector pretraining, ranging from 0 to 100M tokens sampled from FineWeb [33]. The results of this ablation are shown in Figure 4. We found that the optimal value of the $\alpha$ parameter stabilizes around 70M tokens used to pretrain the projectors.
Figure 4: As evaluated on our PTB completions synthetic with Mamba-130M and GPT-Neo-125M, we find that the optimum stabilizes at around 70M tokens of projector pretraining.
@@ Appendix C Additional Pathfinder Results
Appendix D On Baselines
The correct set of baselines for Manticore is an interesting and somewhat challenging question. In the main text, we compare to the set of component models used to construct a Manticore hybrid—in other words, in order for Manticore to be at least as performant as its component models on a task, it must match or beat the performance of the best component model, which implies that both component models need to be fine-tuned. This would roughly match the total amount of fine-tuning FLOPs used to train the corresponding Manticore hybrid. However, there are other potential ways to make a comparison; in this section, we will discuss the fairness and availability of baselines corresponding to different metrics of comparison, and provide a new set of baselines involving ensembles of component models. Specifically, we will address the question of whether the correct comparison is one involving parameter count, training FLOPs, or inference FLOPs.
D.1 Parameter Count
One proposal is to compare a Manticore hybrid of size $N$ to a pretrained model that is also of size $N$. Manticore combines the weights of existing pretrained models to produce a hybrid that is drastically cheaper to generate compared to pretraining a hybrid of the same size from scratch. Off-the-shelf pretrained models of size $N$ are often pretrained up to $D$ tokens corresponding to its Chinchilla optimum [16], but information about the amount, mixture, or quality of pretraining data is often unavailable. This makes comparison along the axis of the parameter count alone somewhat challenging—a larger model may well have been trained on more total data than the two smaller component models making up Manticore. In other words, Manticore should not be expected to follow the same pretraining scaling laws as models that were trained from scratch. Therefore, comparing a Manticore hybrid and a pretrained model of the same size is not necessarily a fair comparison, when considering model size alone. Furthermore, pretrained models of a specific predefined size $N$ are not even guaranteed to exist.
D.2 Training FLOPs
Another option is to make a comparison along the axis of total training FLOPs, which would include pretraining FLOPs, fine-tuning FLOPs, and any additional FLOPs incurred when generating a Manticore hybrid. Suppose we create a Manticore hybrid from two component models of sizes $N_{1}$ and $N_{2}$, which have been pretrained using $T_{1}$ and $T_{2}$ tokens, incurring roughly $6N_{1}T_{1}$ and $6N_{2}T_{2}$ FLOPs, respectively [19]. With Manticore, we incur FLOPs from two sources: projector pretraining and fine-tuning. In our experiments, we use $T_{\text{proj}}=100\text{M}$ tokens of general data for projector pretraining, and saw in Figure 4 that we likely didn’t even need this much. Nonetheless, $100\text{M}$ tokens is substantially smaller than the typical amount of pretraining data, so we can assume that $T_{\text{proj}}=100\text{M}<<\min{\{T_{1},T_{2}\}}$, and since the pretrained projectors can be reused, this cost can be amortized over many future fine-tuning runs. Manticore then involves fine-tuning on some small amount of downstream tasks-specific data comprising $T_{\text{ft}}<<\min{\{T_{1},T_{2}\}}$ tokens. So then, the total amount of training FLOPs involved end-to-end in producing a Manticore hybrid is
$6N_{1}T_{1}+6N_{2}T_{2}+(6N_{1}+6N_{2})T_{\text{proj}}+(6N_{1}+6N_{2})T_{\text{ft}}$ $=O(6N_{1}T_{1}+6N_{2}T_{2}),$
meaning that the total training FLOPs is dominated by the pretraining of the component models. Our experiments in the main text compare Manticore to the better of the two component models, which means that both component models need to be fine-tuned (i.e., the baseline comprises ‘both’ component models). Therefore, if the projector pretraining FLOPs are amortized over many fine-tuning runs, Manticore roughly matches the baseline in terms of training FLOPs. That is, this baseline and Manticore effectively requires $6N_{1}T_{1}+6N_{2}T_{2}+(6N_{1}+6N_{2})T_{\text{ft}}$ FLOPs.
D.3 Inference FLOPs
It is true that our baselines in the main text (which are pairs of component models) are cheaper in terms of inference FLOPs compared to Manticore. In fact, Manticore effectively doubles the inference FLOPs by requiring forward passes through both component models. Here, we include an analysis of inference FLOPs showing that the contribution of the projectors is negligible, and we present an additional baseline—combining the component models into an ensemble that is fine-tuned simultaneously using the same fine-tuning budget as Manticore.
Inference FLOPs analysis.
First, we will compute the general form of the inference FLOPs requirement for a component model. Let $d$ be the embedding dimension, let $t$ be the sequence length, let $L$ be the number of blocks, let $v=|\mathcal{V}|$ be the size of the vocabulary set for our downstream task, and let $B(d,t)$ be the inference FLOPs requirement for the blocks in the component model. Then the inference requirement for a single token prediction from the component model is computed by summing the FLOPs requirements from looking up an embedding, computing forward passes through a sequence of blocks, and generating the final logits. That is, we obtain the following:
$O(1+LB(d,t)+dv)=O(LB(d,t)+dv).$
For a Manticore hybrid, assume that we have $K=2$ component models, $M_{1}$ and $M_{2}$, as well as their projectors. Without loss of generality, assume that the embedding dimensions, $d$, and the number of blocks, $L_{M}$, in the component models are the same. Let $L<<L_{M}$ be the number of Manticore blocks, which is typically constant with respect to the number of blocks in each of the component models $L_{M}$ (in our experiments, $L$ was set to 1 or 2). Let $B_{M_{1}}(d,t)$ and $B_{M_{2}}(d,t)$ be the FLOPs requirements of individual blocks from $M_{1}$ and $M_{2}$ respectively, and let $B_{\text{proj}}(d,t)=O(td^{2})$ be the FLOPs requirement of projector usage. Note that typically, $B_{\text{proj}}(d,t)=O(td^{2})\leq B_{M_{*}}(d,t)$, as many types of blocks involve a dimension-mixing operation such as an MLP, which has a larger FLOPs requirement than $O(td^{2})$, or a sequence mixer that has quadratic or log-linear dependence on $t$, rather than the linear dependence of $B_{\text{proj}}(d)$. Then the FLOPs requirement of each Manticore block is as follows:
$O\left(\frac{L_{M}}{L}(B_{M_{1}}(d,t)+B_{M_{2}}(d,t))+4B_{\text{proj}}(d,t)\right),$
and along with the token embedding and the logits output, we have
$O(1)+L*O\left(\frac{L_{M}}{L}(B_{M_{1}}(d,t)+B_{M_{2}}(d,t))+4tB_{\text{proj}}(d,t)\right)+O(dv)$
$=$ $O\left(L_{M}B_{M_{1}}(d,t)+L_{M}B_{M_{2}}(d,t)+LB_{\text{proj}}(d,t)+dv\right)$
$=$ $O\left(L_{M}B_{M_{1}}(d,t)+L_{M}B_{M_{2}}(d,t)+td^{2}L+dv\right)$
$=$ $O\left(L_{M}B_{M_{1}}(d,t)+L_{M}B_{M_{2}}(d,t)+dv\right),$
where the final step comes from $L<<L_{M}$ and the assumption that $B_{\text{proj}}(d,t)=O(td^{2})\leq B_{M_{*}}(d,t)$. This inference cost is the same as inference with both component models. This motivates another baseline: ensembles of component models, which we evaluate next.
Comparison to ensembles.
We compare the fine-tuning performance of Manticore to ensembles of component models on the six tasks shown in Figure 3. Starting with pretrained Pythia-410M and Mamba-370M models, we construct our ensemble as follows: for each token prediction, we mix the output probabilities from Pythia-410M and Mamba-370M with equal weighting of $0.5$, and then we fine-tune the entire mixture end-to-end on the downstream task. We present the results in Table 8. The ensemble baseline underperforms Manticore and the best component model on all tasks—we suspect that this could be related to overfitting.
Task Pythia-410M Mamba-370M Ensemble Manticore
(A) (B) [A, B] [A, B]
Es. + Alpaca 1.819 1.704 2.172 1.664
Ch. + Alpaca 3.729 3.447 3.854 3.369
Vi. + Alpaca 2.130 2.004 2.173 1.980
NI non-En. 1.764 1.560 1.652 1.530
OpenOrcha 1.570 1.576 1.756 1.553
XQuAD Ar. 0.205 0.207 0.533 0.201
Table 8: Comparison between Manticore, its component models, and an ensemble of its component models on the tasks from Figure 3. For Manticore, we show the best performance achieved across our sweep from Figure 3. Ensembling the component models does not improve performance, but creating a Manticore hybrid does lead to improved performance.
Appendix E Hyperparameters
In this section, we discuss our hyperparameters and our experimental setup. Code implementing our experiments can be found at https://anonymous.4open.science/r/manticore-anon.
E.1 Fine-Tuning Pretrained Hybrids
@@
Penn Treebank completions synthetic. For model weights, we use the AdamW [27] optimizer with a linear learning rate schedule with an initial learning rate of 5e55e-5. For mixture weights, we use the AdamW [27] optimizer with a linear learning rate schedule with an initial learning rate of $0.005$ and use alternating updates.
@@
Fine-tuning on language tasks. For model weights, we use the AdamW [27] optimizer with a linear learning rate schedule with an initial learning rate of 5e55e-5. For mixture weights, we use the AdamW [27] optimizer with a linear learning rate schedule with an initial learning rate of $0.005$ and use simultaneous updates.
E.2 Training Hybrids from Scratch
@@
Existing hybrids were trained with a hyperparameter grid search over the space [1e4,5e4,1e3][1e-4,5e-4,1e-3] for learning rate and $[0.0,0.1]$ for weight decay, similar to the procedure in MAD [37].
Manticore is trained in two stages. In the first stage, we train the model and architecture weights in the alternating schedule utilized in DARTS [26]. In this stage, we perform a hyperparameter grid search of the space [1e4,5e4,1e3][1e-4,5e-4,1e-3] for model weight learning rate, [1e4,1e4][1e-4,1e-4] for architecture weight learning rate, and $[0.1]$ for weight decay. In the second stage, the architecture weights are frozen and we train only the model weights using the best learning rate found in the first stage.
@@
  • ListOps. We trained all models for 5000 steps. GPT-Neo used 8 attention heads, 6 blocks, an embedding dimension of 512, and a feed-forward network (FFN) dimension of 2048. Mamba used 12 blocks with a model dimension of 512. The vocabulary size was 18.
  • IMDb. We trained all models for 25 epochs with a batch size of 32. GPT-Neo used 8 attention heads, 6 blocks, an embedding dimension of 512, and an FFN dimension of 2048. Mamba used 12 blocks with a model dimension of 512. The vocabulary size was 129.
  • CIFAR10. We trained all models for 10 epochs. GPT-Neo used 4 attention heads, 3 blocks, an embedding dimension of 64, and an FFN dimension of 128. Mamba used 6 blocks with a model dimension of 64. The vocabulary size was 256, corresponding to the pixel value range of the grayscale image.
  • Pathfinder32. We trained all models for 10 epochs. GPT-Neo used 8 attention heads, 4 blocks, an embedding dimension of 128, and an FFN dimension of 128. Mamba used 8 blocks with a model dimension of 128. The vocabulary size was 256, corresponding to the pixel value range of the grayscale image.
@@
We use two blocks each from GPT-Neo and Mamba, each with a model dimension of 128. We train for 200 epochs and select the best performance during training, as all of the models overfit across the board. We use the AdamW [27] optimizer with a linear learning rate schedule with an initial learning rate of 5e55e-5.
E.3 Programming Hybrids
@@
MAD tasks as a search proxy. For model weights, we use the AdamW [27] optimizer with a linear learning rate schedule with an initial learning rate of 5e55e-5. For mixture weights, we use the AdamW [27] optimizer with a linear learning rate schedule with an initial learning rate of $0.01$ and use simultaneous updates. For search on the MAD tasks, we train scaled-down versions of GPT-Neo and Mamba each with four blocks, model dimensions of 128, and no projectors.
E.4 Pretraining Projectors
For all non-frozen weights (i.e., projectors, mixture weights, embeddings, and the LM head), we use the AdamW [27] optimizer with a linear learning rate schedule with an initial learning rate of 5e55e-5.
Appendix F Data and MAD Task Parameters
@@
Natural language tasks. We evaluate the ability to fine-tune Manticore on natural language datasets. Specifically, we evaluate on Penn Treebank [29], the Alpaca instruction tuning dataset [40], and an i.i.d. split of the ELI5 training set [10]. Additionally, we use 100M tokens from the FineWeb dataset [33] to pretrain our projector weights. We describe all other natural language datasets that we use in our evaluations below.
  • NI Spanish QA + Alpaca. This is from the Natural Instruction dataset v2.8 downloaded from https://github.com/allenai/natural-instructions/releases, we picked task 1610 and mixed it with equal numbers of randomly selected samples from the Alpaca dataset to create a bilingual dataset that contains Spanish Q&A along with English instructions.
  • NI Chinese QA + Alpaca. This is similar to the previous dataset, except we pick task1570, which is Q&A that input/output language are Chinese.
  • MLQA Vietnamese + Alpaca. This dataset is a subset of MLQA (MultiLingual Question Answering)(https://huggingface.co/datasets/facebook/mlqa) in which both the inputs and outputs are in Vietnamese, and mixed with equal numbers of randomly selected samples from Alpaca dataset to create a bilingual dataset.
  • OpenOrcha. We randomly sample 10,000 samples from the OpenOrcha dataset containing Japanese translations from https://huggingface.co/datasets/atsushi3110/cross-lingual-openorcha-830k-en-ja, to form a Japanese Q&A dataset.
  • NI all non-English QA. There are six Q&A tasks in the Natural Instructions dataset such that both their input and output language is non-English—we combine all of them to form a new dataset containing non-English Q&A.
  • XQuAD Arabic. The Arabic Q&A part from XQuAD (Cross-lingual Question Answering Dataset), from https://huggingface.co/datasets/google/xquad.
Appendix G A Call for Action & Community Recommendations
@@
We found that standard tools such as Huggingface and PyTorch were insufficient to cleanly access intermediate activations across several model implementations. This could be resolved by adopting standard implementations or structures for LMs that share the common block structure that we describe in Section 3.1. Instead, our solution was to fork implementations of several Huggingface models, which is time-consuming, error-prone, and non-scalable. A solution to this problem would enable and encourage further research on pretrained hybrid models, which in turn helps to democratize LM research.
@@ Appendix H Limitations
Appendix I Compute Resources